8C4/12C4 perform the indicated calculationn
Answer by Theo(13342)

if you have a calculator that does this for you, you will get.....

8C4 = 70
12C4 = 495
8C4 / 12C4 = 70/495
this simplifies to 14/99.

if you don't have a calculator that does this for you, then you have to do it by formula.

the combination formula is nCx = n! / (x! * (n-x)!)

you get:

8C4 = 8! / (4! * 4!)

this is the same as:

8C4 = (8 * 7 * 6 * 5 * 4!) / (4! * 4!)

the 4! in the numerator and denominator cancel out and you are left with:

8C4 = (8 * 7 * 6 * 5) / (4 * 3 * 2 * 1)

solve for 8C4 to get:

8C4 = 70

and you get:

12C4 = 12! / (4! * 8!)

this is the same as:

12C4 = (12 * 11 * 10 * 9 * 8!) / (4! * 8!)

the 8! in the numerator and denominator cancel out and you are left with:

12C4 = (12 * 11 * 10 * 9) / (4 * 3 * 2 * 1)

solve for 12C4 to get:

12C4 = 495

8C4 / 12C4 = 70 / 495 = 14 / 99, same as before.

in decimal format, it is equal to .141414......